Ebony Reigns owns a studio that would cost ¢ 120,000 to replace should it ever be destroyed by fire. There is a 25% chance that the studio could be destroyed by fire during the course of the year. If the fire occurs, Ebony Reign's studio will be worth only ¢ 60,000. An insurance company has offered Ebony a false insurance policy that requires her to pay a yearly premium of ¢ 15,000 in the good state of nature (no fire) Ebony has fully insured her studio to eliminate the risk. Assuming that Ebony Reigns is risk averse has another wealth answer the following questions:



e) Calculate the variance of the value of Ebony's studio with fair insurance. [4marks]



f) Is Ebony better off with the fair insurance ? Why ? [4marks]

1
Expert's answer
2021-11-16T11:35:24-0500

(a)

100%=120,000

75%"=\frac{120,000\times 75}{100}=90,000"


(b)

Without insurance, Ebony faces a risk of losing:

"120,000-60,000=60,000."


(c)

Under fair insurance, in bad state of nature, Ebony should be paid: 120,000 which is the total worth of her property.


(d)

Ebony is risk averse, she has another wealth in case there is no insurance.

With insurance, all the total worth of her property will be compensated for.